OpenStudy (anonymous):

@YanaSidlinskiy can u help with this promblem Isadora bought 4 trees to plant in her backyard and spent more than one hundred twenty-eight dollars altogether. How many dollars could each tree have cost? Let t represent how many dollars each tree cost.

OpenStudy (yanasidlinskiy):

Why did you close it?? I was going to help you..but couldn't cuz of the site. I kept on loosing connection..Mind re-opening it??

OpenStudy (yanasidlinskiy):

This question has to deal with the inequality. The inequality here is: \[4t \ge 84\] the question is: how much could EACH tree cost? meaning you will have to divide 4t by 4 and also 84 by 4 this gives you: \[\frac{ 4t }{ 4 } \ge \frac{ 84 }{ 4 }\]

OpenStudy (yanasidlinskiy):

What do you get?